# A Synthetic Benchmark for Collaborative 3D Semantic Occupancy Prediction in V2X Autonomous Driving

### v1

@@ Abstract
3D semantic occupancy prediction is an emerging perception paradigm in autonomous driving, providing a voxel-level representation of both geometric details and semantic categories. However, the perception capability of a single vehicle is inherently constrained by occlusion, restricted sensor range, and narrow viewpoints. To address these limitations, collaborative perception enables the exchange of complementary information, thereby enhancing the completeness and accuracy. In the absence of a dedicated dataset for collaborative 3D semantic occupancy prediction, we augment an existing collaborative perception dataset by replaying it in CARLA with a high-resolution semantic voxel sensor to provide dense and comprehensive occupancy annotations. In addition, we establish benchmarks with varying prediction ranges designed to systematically assess the impact of spatial extent on collaborative prediction. We further develop a baseline model that performs inter-agent feature fusion via spatial alignment and attention aggregation. Experimental results demonstrate that our baseline model consistently outperforms single-agent models, with increasing gains observed as the prediction range expands.
Introduction
@@
Collaborative perception has emerged as a powerful strategy to enhance scene understanding in autonomous driving (Liu et al. 2023; Huang et al. 2023). By sharing sensory information across multiple agents, it enables vehicles to perceive a broader environment beyond their individual field of view. This is particularly crucial in urban driving scenarios where occlusion and limited sensor range frequently hinder the perception performance of single vehicles. Prior research (Xu et al. 2022a; Hu et al. 2022; Yang et al. 2024) has shown that inter-agent communication and feature fusion can significantly improve the accuracy of 3D detection, segmentation, and tracking tasks.
@@
However, most existing collaborative perception methods employ coarse environment representations, such as 3D bounding boxes or bird’s-eye-view (BEV) maps. While effective for certain tasks, these representations are insufficient for capturing the fine-grained geometry and semantics necessary for downstream reasoning and planning. Recently, 3D semantic occupancy prediction, also known as Semantic Scene Completion (SSC), has gained attention for its ability to provide detailed 3D semantic and geometric information by utilizing a fine-grain voxel-based representation (Xu et al. 2024; Shi et al. 2023; Zhang et al. 2024b), offering a richer and more detailed understanding of the environment.
@@
Despite its potential, 3D semantic occupancy prediction remains largely unexplored in collaborative settings. This is due in part to the absence of dedicated datasets and benchmarks that provide voxel-level semantic ground-truth annotations for multiple agents under collaborative settings (Teufel et al. 2024; Yazgan, Akkanapragada, and Zöllner 2024). Existing datasets, such as SemanticKITTI (Behley et al. 2019) and Occ3D (Tian et al. 2024), lack support for multi-agent configurations and primarily rely on LiDAR point cloud of single vehicle to generate annotations. However, the inherent sparsity of LiDAR point clouds, particularly at greater distances, coupled with occlusions, sensor noise, and non-uniform point distributions, poses significant challenges to generating accurate and reliable annotations. Moreover, collaborative perception imposes higher demands on annotations, requiring temporally and spatially aligned multi-agent labels to ensure consistent supervision and effective fusion.
@@
To address this gap, we augment an existing multi-agent dataset (Xu et al. 2022b) by replaying it in CARLA with a high-resolution semantic voxel sensor, yielding dense voxel-wise annotations for both geometry and semantics. The resulting dataset, which we term Co3SOP, supports the training and evaluation of collaborative 3D semantic occupancy prediction models under simulation scenarios. Alongside the dataset, we establish benchmarks with different prediction ranges, to systematically assess how spatial distance influences the effectiveness of collaboration. Additionally, we introduce a baseline model that incorporates inter-agent feature fusion with spatial alignment and sparse attention. Specifically, we apply warping-based spatial alignment to transform neighboring agents’ features into the ego frame, followed by a visibility-guided sparse attention mechanism modulated by a learned confidence mask to adaptively weight contributions from each agent.
@@
Table 1: Comparison of the existing 3D Occupancy Prediction Dataset for Autonomous Driving and our proposed Co3SOP. Real and Sim represent the raw data is collected from real world or simulation platform. Gen means the 3D occupancy annotations are generated from raw data. C, L denote camera and LiDAR.
@@
SSCBench Real+Gen nuScenes&Waymo &KITTI-360 C&L [256, 256, 32] [0.2, 0.2, 0.2] -
V2VSSC Sim+Gen OPV2V C+L [128, 128, 20] [0.78, 0.78, 0.4] V2V
Co3SOP Sim OPV2V C+L [1000, 1000, 70] [0.1, 0.1, 0.1] V2V
@@
Related Work
3D Semantic Occupancy Prediction
Methodologies: 3D semantic occupancy prediction aims to provide voxel-level understanding of both geometry and semantics, and has seen rapid development across different sensing modalities. Early approaches such as SSCNet (Song et al. 2017) adopt volumetric CNNs on RGB-D inputs. LiDAR-based methods like LMSCNet (Roldão, de Charette, and Verroust-Blondet 2020), S3CNet (Cheng et al. 2021), and JS3CNet (Yan et al. 2021) exploit the structural sparsity and accuracy of 3D point clouds to produce high-quality voxel predictions. Meanwhile, vision-based methods, including MonoScene (Cao and de Charette 2022), infer 3D semantics by lifting monocular depth and segmentation into voxel space. To overcome the limitations of single-view inputs, recent works such as VoxFormer (Li et al. 2023b) and OccFormer (Zhang, Zhu, and Du 2023) leverage multi-camera fusion and transformer-based 2D-to-3D attention mechanisms. Despite steady progress, these methods are restricted to single-agent perception, and thus remain vulnerable to occlusion, field-of-view limitations, and sensor sparsity.
Datasets: To facilitate the development of 3D semantic occupancy prediction, several datasets have been proposed with voxel-level annotations derived primarily from LiDAR scans. SemanticKITTI (Behley et al. 2019, 2021; Geiger, Lenz, and Urtasun 2012) is a widely adopted benchmark, providing dense semantic labels on voxelized point clouds. KITTI-360 (Liao, Xie, and Geiger 2022) extends this setup with panoramic imagery and 360° LiDAR coverage. Occ3D (Tian et al. 2024) introduces a voxel annotation pipeline based on Waymo and nuScenes data, offering large-scale semantic occupancy labels. SSCBench (Li et al. 2023a) further consolidates multiple sources to unify evaluation across diverse urban scenes. While these datasets advance single-agent occupancy prediction, they do not support multi-agent collaboration. V2VSSC (Zhang et al. 2024a) is the first to target vehicle-to-vehicle (V2V) collaborative occupancy tasks, yet its annotations remain limited by the sparsity and occlusion inherent in LiDAR data, compromising label density and accuracy. In contrast, our proposed Co3SOP dataset offers simulation-based dense annotations specifically tailored for collaborative settings, as summarized in table 1.
@@
Collaborative Perception
Methodologies: Collaborative perception enhances environmental understanding by enabling vehicles to exchange complementary information. A variety of methods have been proposed to realize effective multi-agent fusion. V2VNet (Wang et al. 2020) adopts graph neural networks to aggregate features from spatially distributed agents. V2X-ViT (Xu et al. 2022a) leverages vision transformers to model inter-agent interactions through attention mechanisms. To improve communication efficiency, Where2comm (Hu et al. 2022) identifies task-relevant regions for selective transmission, while How2comm (Yang et al. 2024) explores policies for adaptive communication scheduling. While effective for tasks such as object detection and segmentation, these approaches are limited to object-centric outputs and fall short in supporting fine-grain voxel-level semantic understanding.
Datasets: Ego-vehicle datasets such as KITTI (Mei et al. 2022), nuScenes (Caesar et al. 2020), and Waymo (Mei et al. 2022) fall short in supporting collaborative perception, as they lack multi-agent scenarios and aligned multi-view annotations. To address these limitations, a range of specialized datasets have emerged (Teufel et al. 2024; Yazgan, Akkanapragada, and Zöllner 2024; Zimmer et al. 2024; Hao et al. 2024). Among simulated datasets, V2X-Sim (Li et al. 2022) models V2X interactions in CARLA and supports tasks such as detection, tracking, and BEV segmentation. OPV2V (Xu et al. 2022b), also built in CARLA, emphasizes real-time V2V communication and provides configurable multi-agent scenarios. In real-world settings, V2V4Real (Xu et al. 2023) focuses on vehicle-to-vehicle collaboration, while DAIR-V2X (Yu et al. 2022) targets vehicle-to-infrastructure (V2I) scenarios. Additionally, V2X-Seq (Yu et al. 2023) supports sequential V2X perception by enabling temporal information sharing across agents. However, none of the above datasets provide dense voxel-level semantic annotations required for collaborative 3D semantic occupancy prediction.
Figure 2: Illustration of the V2V scenarios in Carla and the corresponding data collection results. Left: The screen shot of three V2V scenarios in Carla based on the settings in OPV2V. Mid: LiDAR generated 3D semantic voxel annotations. Right: The annotations collected by our developed 3D semantic voxel sensor.
Co3SOP Dataset
Semantic Voxel Annotation Pipeline
We construct the Co3SOP dataset by replaying multi-agent driving scenarios from OPV2V (Xu et al. 2022b) in the CARLA simulator and augmenting them with dense voxel-level semantic occupancy labels. Unlike prior datasets that rely on sparse LiDAR observations, Co3SOP leverages CARLA’s high-fidelity simulation environment to generate complete ground truth.
@@
However, dividing a large scene into high-resolution voxels and checking each voxel individually, however, can introduce significant computational overhead. For example, a detection range of $100\times 100\times 4.8\,m^{3}$ at $0.1\,m$ resolution results in approximately $48$ million detection operations for one frame. To alleviate this, we design a multi-stage voxel annotation pipeline, as illustrated in fig. 1 that leverages scene-level object cues to minimize redundant computations:
@@
Object-Level Occupancy Completion. Then, for each object, a parallelized Breadth-First Search (BFS) is performed starting from its seed voxels. The BFS selectively expands to the six neighboring voxels (top, bottom, left, right, front, back), performing fine-grained collision checks to determine occupancy. If a neighbor is found to be occupied, it is added to the seed voxels. This localized exploration continues iteratively until all reachable occupied voxels are annotated, effectively reconstructing the complete volumetric occupancy of the object.
Object Label Assignment. After occupancy completion, the semantic category of the intersecting object—retrieved from the CARLA simulation engine—is assigned to each of its occupied voxels, completing the voxel-wise annotation process.
@@
To validate the accuracy and completeness of our annotations, we conduct a visual comparison against labels generated from LiDAR scans using the annotation way in SurroundOcc (Wei et al. 2023). Figure 2 illustrates representative V2V scenarios, highlighting voxel labels produced by both approaches. Figure 2 illustrates representative V2V scenarios, highlighting voxel labels produced by both approaches. As illustrated, our pipeline yields significantly denser and more complete semantic voxel annotations, especially in occluded or long-range regions where LiDAR-based annotations suffer from sparsity and missing data.
Co3SOP inherits the multi-agent traffic scenarios of OPV2V (Xu et al. 2022b), retaining its training, validation, and testing splits. Distinctively, we treat each vehicle in a scene as an independent training target, allowing every agent to serve as the ego vehicle while dynamically collaborating with its surrounding vehicles. This formulation not only maximizes data utilization, but also enables flexible many-to-many collaboration configurations across agents.
@@
The overall structure of Co3SOP-Base is illustrated in fig. 3. the framework first uses a shared backbone network to extract multi-scale image features from each view independently. These 2D features are then lifted into the 3D voxel space via an image deformable cross-attention module. Once voxel features are obtained for each vehicle, collaborative fusion is enabled by sharing features among connected vehicles (CVs). Upon receiving voxel features from neighboring vehicles, the ego vehicle performs 3D affine transformation to align these features into its own coordinate frame, ensuring proper spatial alignment. Then, the ego vehicle fuses the aligned features with its own voxel representation using a voxel deformable cross-attention mechanism, modulated by a hybrid attention mask that integrates warping and confidence priors. Finally, a prediction head with 3D convolution predicts the voxel-level semantic occupancy map. The details of the key modules are as follows.

### v2

# A Synthetic Benchmark for Collaborative 3D Semantic Occupancy Prediction in V2X-Enabled Autonomous Driving

@@ Abstract
3D semantic occupancy prediction is an emerging perception paradigm in autonomous driving, providing a voxel-level representation of both geometric details and semantic categories. However, its effectiveness is inherently constrained in single-vehicle setups by occlusions, restricted sensor range, and narrow viewpoints. To address these limitations, collaborative perception enables the exchange of complementary information, thereby enhancing the completeness and accuracy of predictions. Despite its potential, research on collaborative 3D semantic occupancy prediction is hindered by the lack of dedicated datasets. To bridge this gap, we design a high-resolution semantic voxel sensor in CARLA to produce dense and comprehensive annotations. We further develop a baseline model that performs inter-agent feature fusion via spatial alignment and attention aggregation. In addition, we establish benchmarks with varying prediction ranges designed to systematically assess the impact of spatial extent on collaborative prediction. Experimental results demonstrate the superior performance of our baseline, with increasing gains observed as range expands.
I Introduction
@@
Collaborative perception has emerged as a powerful strategy to enhance scene understanding in autonomous driving [1, 2]. By sharing sensory information across multiple agents, it enables vehicles to perceive a broader environment beyond their individual field of view. This is particularly crucial in urban driving scenarios where occlusions and limited sensor range frequently hinder the perception performance of a single vehicle. Prior research [3, 4, 5] has shown that inter-agent communication and feature fusion can significantly improve the accuracy of 3D detection, segmentation, and tracking tasks.
@@
However, most existing collaborative perception methods employ coarse environment representations, such as 3D bounding boxes or bird’s-eye-view (BEV) maps. While effective for certain tasks, these representations are insufficient for capturing the fine-grained geometry and semantics necessary for downstream reasoning and planning. Recently, 3D semantic occupancy prediction, also known as Semantic Scene Completion (SSC), has gained attention for its ability to provide detailed 3D semantic and geometric information by utilizing a fine-grain voxel-based representation [6, 7, 8], offering a richer and more detailed understanding of the environment.
@@
Despite its potential, 3D semantic occupancy prediction remains largely unexplored in collaborative settings. This is due in part to the absence of dedicated datasets and benchmarks that provide voxel-level semantic ground-truth annotations under collaborative settings [9, 10]. Existing datasets, such as SemanticKITTI [11] and Occ3D [12], lack support for multi-agent configurations and primarily rely on LiDAR point cloud of single vehicle to generate annotations. However, the inherent sparsity of LiDAR point clouds, particularly at greater distances, coupled with occlusions, sensor noise, and non-uniform point distributions, poses significant challenges to generating accurate and reliable annotations. Moreover, collaborative perception imposes higher demands on annotations, requiring temporally and spatially aligned multi-agent labels to ensure consistent supervision and effective fusion.
@@
To address this gap, we augment an existing multi-agent dataset [13] by replaying it in CARLA with a high-resolution semantic voxel sensor, yielding dense voxel-level annotations. The resulting dataset, which we term Co3SOP, supports the training and evaluation of collaborative 3D semantic occupancy prediction models under simulation scenarios. Additionally, we introduce a baseline model that incorporates inter-agent feature fusion with spatial alignment and sparse attention. Specifically, we apply warping-based spatial alignment to transform neighboring agents’ features into the ego frame, followed by a visibility-guided sparse attention mechanism modulated by a learned confidence mask to adaptively weight contributions from each agent. Alongside the dataset and baseline, we establish benchmarks with different prediction ranges, to systematically assess how spatial distance influences the effectiveness of collaboration.
@@
TABLE I: Comparison of the existing 3D Occupancy Prediction Dataset for Autonomous Driving and our proposed Co3SOP. Real and Sim represent the raw data is collected from real world or simulation platform. Gen means the 3D occupancy annotations are generated from raw data. C, L denote camera and LiDAR.
@@
SSCBench Real+Gen
nuScenes&Waymo
&KITTI-360
C&L [256, 256, 32] [0.2, 0.2, 0.2] -
V2VSSC Sim+Gen OPV2V C&L [128, 128, 20] [0.78, 0.78, 0.4] V2V
Co3SOP (Ours) Sim OPV2V C&L [1000, 1000, 70] [0.1, 0.1, 0.1] V2V
@@
II Related Work
II-A 3D Semantic Occupancy Prediction
Methodologies: 3D semantic occupancy prediction aims to provide voxel-level understanding of both geometry and semantics, and has seen rapid development across different sensing modalities. Early approaches such as SSCNet [14] adopt volumetric CNNs on RGB-D inputs. LiDAR-based methods like LMSCNet [15], S3CNet [16], and JS3CNet [17] exploit the structural sparsity and accuracy of 3D point clouds to produce high-quality voxel predictions. Meanwhile, vision-based methods, including MonoScene [18], infer 3D semantics by lifting monocular depth and segmentation into voxel space. To overcome the limitations of single-view inputs, recent works such as VoxFormer [19] and OccFormer [20] leverage multi-camera fusion and transformer-based 2D-to-3D attention mechanisms. Despite steady progress, these methods are restricted to single-agent perception, and thus remain vulnerable to occlusion, field-of-view limitations, and sensor sparsity.
Datasets: To facilitate the development of 3D semantic occupancy prediction, several datasets have been proposed with voxel-level annotations derived primarily from LiDAR scans. SemanticKITTI [11, 21, 22] is a widely adopted benchmark, providing dense semantic labels on voxelized point clouds. KITTI-360 [23] extends this setup with panoramic imagery and 360° LiDAR coverage. Occ3D [12] introduces a voxel annotation pipeline based on Waymo and nuScenes data, offering large-scale semantic occupancy labels. SSCBench [24] further consolidates multiple sources to unify evaluation across diverse urban scenes. While these datasets advance single-agent occupancy prediction, they do not support multi-agent collaboration. V2VSSC [25] is the first to target vehicle-to-vehicle (V2V) collaborative occupancy tasks, yet its annotations remain limited by the sparsity and occlusion inherent in LiDAR data, compromising label density and accuracy. In contrast, our proposed Co3SOP dataset offers simulation-based dense annotations specifically tailored for collaborative settings, as summarized in table I.
@@ II Related Work
II-B Collaborative Perception
Methodologies: Collaborative perception enhances environmental understanding by enabling vehicles to exchange complementary information. A variety of methods have been proposed to realize effective multi-agent fusion. V2VNet [26] adopts graph neural networks to aggregate features from spatially distributed agents. V2X-ViT [3] leverages vision transformers to model inter-agent interactions through attention mechanisms. To improve communication efficiency, Where2comm [4] identifies task-relevant regions for selective transmission, while How2comm [5] explores policies for adaptive communication scheduling. While effective for tasks such as object detection and segmentation, these approaches are limited to object-centric outputs and fall short in supporting fine-grain voxel-level semantic understanding.
Datasets: Ego-vehicle datasets such as KITTI [27], nuScenes [28], and Waymo [27] fall short in supporting collaborative perception, as they lack multi-agent scenarios and aligned multi-view annotations. To address these limitations, a range of specialized datasets have emerged [9, 10, 29, 30]. Among simulated datasets, V2X-Sim [31] models V2X interactions in CARLA and supports tasks such as detection, tracking, and BEV segmentation. OPV2V [13], also built in CARLA, emphasizes real-time V2V communication and provides configurable multi-agent scenarios. In real-world settings, V2V4Real [32] focuses on vehicle-to-vehicle collaboration, while DAIR-V2X [33] targets vehicle-to-infrastructure (V2I) scenarios. Additionally, V2X-Seq [34] supports sequential V2X perception by enabling temporal information sharing across agents. However, none of the above datasets provide dense voxel-level semantic annotations required for collaborative 3D semantic occupancy prediction.
III Co3SOP Dataset
III-A Semantic Voxel Annotation Pipeline
We construct the Co3SOP dataset by replaying multi-agent driving scenarios from OPV2V [13] in the CARLA simulator and augmenting them with dense voxel-level semantic occupancy labels. Unlike prior datasets that rely on sparse LiDAR observations, Co3SOP leverages CARLA’s high-fidelity simulation environment to generate complete ground truth.
@@ III-A Semantic Voxel Annotation Pipeline
Dividing a large scene into high-resolution voxels and checking each voxel individually, however, can introduce significant computational overhead. For example, a detection range of $100\times 100\times 4.8\,m^{3}$ at 0.1m0.1\,m resolution results in approximately $48$ million detection operations for one frame. To alleviate this, we design a multi-stage voxel annotation pipeline, as illustrated in fig. 1 that leverages scene-level object cues to minimize redundant computations:
@@
Object-Level Occupancy Completion. Then, for each object, a parallel Breadth-First Search (BFS) is initiated from its seed voxels. The BFS selectively propagates to the six axis-aligned neighboring voxels (i.e., $\pm x,\pm y$, and $\pm z$), where each candidate undergoes a collision-aware check for occupancy. Occupied occupied voxels are added to the expansion frontier for subsequent iterations. This localized and iterative process continues until all reachable occupied voxels are exhaustively annotated. The entire design is inherently multi-threaded, with each object assigned to an independent thread, ensuring efficient and scalable occupancy annotation.
Figure 2: Illustration of the V2V scenarios in Carla and the corresponding data collection results. Left: The screen shot of two V2V scenarios in Carla based on the settings in OPV2V. Mid: LiDAR generated 3D semantic voxel annotations. Right: The annotations collected by our developed 3D semantic voxel sensor.
Object Label Assignment. After occupancy completion, the semantic category of the intersecting object—retrieved from the CARLA simulation engine—is assigned to each of its occupied voxels, completing the voxel-level annotation process.
@@ III-B Annotation Statistics
To validate the accuracy and completeness of our annotations, we conduct a visual comparison against labels generated from LiDAR scans using the annotation way in SurroundOcc [35]. Figure 2 illustrates representative V2V scenarios, highlighting voxel labels produced by both approaches. Figure 2 illustrates representative V2V scenarios, highlighting voxel labels produced by both approaches. As illustrated, our pipeline yields significantly denser and more complete semantic voxel annotations, especially in occluded or long-range regions where LiDAR-based annotations suffer from sparsity and missing data.
Co3SOP inherits the multi-agent traffic scenarios of OPV2V [13], retaining its training, validation, and testing splits. Distinctively, we treat each vehicle in a scene as an independent training target, allowing every agent to serve as the ego vehicle while dynamically collaborating with its surrounding vehicles. This formulation not only maximizes data utilization, but also enables flexible many-to-many collaboration configurations across agents.
@@ IV-A Overall Structure
The overall structure of Co3SOP-Base is illustrated in fig. 3. The framework first uses a shared backbone network to extract multi-scale image features $\{X_{i}^{l}\}_{l=1}^{L},X_{i}^{l}\in R^{N\times H\times W\times C}$ for each vehicle $V_{i}$, where $L$ represent the scales amount of image features and $i$ is vehicle id. These 2D features are then lifted into the 3D voxel space via an image deformable cross-attention module. Once voxel features $F_{i}\in R^{X\times Y\times Z\times C}$ are obtained for each vehicle, collaborative fusion is enabled by sharing features among connected vehicles (CVs). Upon receiving voxel features from neighboring vehicles, the ego vehicle performs 3D affine transformation to align these features into its own coordinate frame, ensuring proper spatial alignment. Then, the ego vehicle fuses the aligned features with its own voxel representation using a voxel deformable cross-attention mechanism, modulated by a hybrid attention mask that integrates warping and confidence priors. Finally, a prediction head with 3D convolution predicts the voxel-level semantic occupancy map. The details of the key modules are as follows.
Image Backbone. We adopt ResNet101-DCN [36] as the image backbone to extract multi-scale image features from multi-view camera images. The extracted features are further refined using a Feature Pyramid Network (FPN) [37] to aggregate contextual information across scales.
Image To Voxel Transformation. To integrate the 2D image feature into the 3D space, we apply a image deformable cross-attention as in SurroundOcc [35]. This image deformable cross-attention mechanism aggregates multi-view 2D features and learns to map them into the corresponding 3D voxel locations, by considering geometric relationships, ensuring that spatial consistency is maintained when transitioning from 2D to 3D space:
$MSDeformAttn(q,p,x)=$
$\sum\limits_{m=1}^{N_{head}}W_{m}[\sum\limits_{l=1}^{L}\sum\limits_{k=1}^{K}A_{mlk}$ $\cdot W_{m}^{{}^{\prime}}x_{i}^{l}(\phi_{l}(p_{q})+\Delta p_{qmlk})]$
$f_{i}=MSDeformAttn(q_{i},p_{q}$ $,\{x_{i}^{l}\}_{l=1}^{L})$ (1)
where $q_{i}\in Q_{i}$ is the corresponding position of query, $Q_{i}$ is a set of learnable voxel query, $W_{m}$ and $W_{m}^{{}^{\prime}}$ are learnable weights, $A_{mlk}$ is attention weight, $p_{q}$ is the reference point of the voxel on the image and $\phi_{l}$ indicate the rescale sampling function.
Collaborative Vehicle Spatial Alignment. Before multi-agent fusion, voxel features from neighboring agents are transformed into the ego vehicle’s coordinate frame using a 3D affine transformation module inspired by Spatial Transformer Networks [38]. The transformation matrix is computed from relative poses between agents, ensuring that voxel features from different coordinates are properly aligned for subsequent attention-based fusion:
$F_{j}^{{}^{\prime}}=AffineGrid(F_{j},T_{ji}^{4\times 4})$ (2)
where $AffineGrid$ is a projective transformation function and $T_{ji}^{4\times 4}$ is the pose transformation matrix from vehicle $j$ to vehicle $i$.
Confidence Masked Feature Fusion. To enable robust and adaptive fusion, each vehicle maintains a learnable confidence mask that estimates the voxel-level reliability of its features. In addition, a binary warping mask is generated during feature alignment to indicate valid projection areas. These two masks are combined to form a hybrid attention mask, which modulates the voxel deformable cross-attention. This design allows the ego vehicle to selectively aggregate features from its collaborators based on both alignment accuracy and confidence, suppressing unreliable or noisy regions. The voxel deformable attention dynamically samples neighboring points across agents, guided by learned offsets and the hybrid attention mask:
$f_{i}^{{}^{\prime}}=MSDeformAttn(f_{i},p_{ij},\{f_{j}\}_{j=1}^{J})$ (3)
where $p_{ij}$ is the reference point that $\{f_{j}\}_{j=1}^{J}$ has higher confidence than $f_{i}$.
IV-B Loss
To supervise collaborative 3D occupancy prediction, we combine a voxel-level cross-entropy loss for semantic classification, a scene-class affinity loss [18] to encourage consistency between semantic and geometric features, and a confidence loss that regularizes the learned confidence mask for inter-agent feature consistency.
V Experiments
V-A Dataset and Metrics
To evaluate the performance of collaborative 3D semantic occupancy prediction under varying spatial conditions, we define three perception range settings in our benchmark: $25.6\times 25.6\times 4.8\,m^{3}$ range with 0.1m0.1\,m voxel size, $51.2\times 51.2\times 4.8\,m^{3}$ range with 0.2m0.2\,m voxel size, and $76.8\times 76.8\times 4.8\,m^{3}$ range with 0.3m0.3\,m voxel size. These settings allow us to assess how collaborative perception improves scene understanding as the coverage area increases, particularly in occluded or distant regions where a single vehicle’s sensing capability is limited. For evaluation, we adopt Intersection-over-Union (mIoU), including IoU for each class individually, as well as the mean IoU (mIoU) across all semantic categories.